# BFA: Best-Feature-Aware Fusion for Multi-View Fine-grained Manipulation

### v2

@@ V CONCLUSIONS
In this paper, we propose the Best-Feature-Aware fusion strategy for multi-view fine-grained manipulation. Such strategy achieves the dynamic view fusion during manipulation. It greatly reduces the visual information redundancy and computational costs while significantly improving the success rate of complex fine-grained manipulation tasks. To implement the strategy, we further introduce the VLM-based scoring system to generate the multi-view ground-truth of importance score, as the supervision of the introduced light-weight scoring network. The proposed BFA strategy provides 22%-46% improvements on fine-grained manipulation tasks. In the future, we will extend our method to VLA (Vision-Language-Action) based approaches [28, 29, 30, 31, 32]. Since most tokens of VLA in these method are image tokens, we anticipate our method’s effectiveness and its potential to significantly reduce large computational resources.

### v3

@@ V CONCLUSIONS
In this paper, we propose the Best-Feature-Aware fusion strategy for multi-view fine-grained manipulation. Such strategy achieves the dynamic view fusion during manipulation. It greatly reduces the visual information redundancy and computational costs while significantly improving the success rate of complex fine-grained manipulation tasks. To implement the strategy, we further introduce the VLM-based scoring system to generate the multi-view ground-truth of importance score, as the supervision of the introduced light-weight scoring network. The proposed BFA strategy provides 22%-46% improvements on fine-grained manipulation tasks. In the future, we will extend our method to VLA (Vision-Language-Action) based approaches [27, 32, 33]. Since most tokens of VLA in these method are image tokens, we anticipate our method’s effectiveness and its potential to significantly reduce large computational resources.